OpenStudy (sen_turtle):

Please help Simplify: Sqrt(20/3n^3)

OpenStudy (ddcamp):

What are you trying to find?

OpenStudy (sen_turtle):

what the simplified version is

OpenStudy (ddcamp):

I'm not sure that can be simplified. Can you try typing it with the equation editor?

OpenStudy (sen_turtle):

\[\sqrt{20/3n^3}\]

OpenStudy (photon336):

I'm wondering if we can re-write it like this and then simplify \[\sqrt{\frac{ 20 }{ 3n^{3 }}} = \sqrt{20}*\sqrt{\frac{ 1 }{ 3n^3 }}\]

OpenStudy (erak):

I get something funny like \[\frac{ 2\sqrt{15n ^{3}} }{ 3n ^{3} }\]

OpenStudy (sen_turtle):

I believe the answer is something like that

OpenStudy (erak):

oh but that's only if you're required to rationalize, if not it's just 2sqrt(5)/sqrt(3n^3)

OpenStudy (sen_turtle):

how do you get the 15 on the top?

OpenStudy (erak):

when you factor sqrt(20), it becomes sqrt(4*5), and it can be written as 2sqrt(5). When you have the sqrt(3n^3) on the bottom, you usually multiply fraction by sqrt(3n^3)/sqrt(3n^3) to get rid of the sqrt on the bottom, so the top is 2sqrt(5)*sqrt(3n^3), and the 5 goes into the sqrt(3n^3), making it sqrt(15n^3) Also i'm sorry the equation maker thing takes me too long

OpenStudy (sen_turtle):

okay thank you

jimthompson5910 (jim_thompson5910):

See the attached PDF for a similar example.

OpenStudy (erak):

@jim_thompson5910 Do you agree with my answer? Just wondering

jimthompson5910 (jim_thompson5910):

Which answer @Erak this one you wrote here? \[\Large \frac{ 2\sqrt{15n ^{3}} }{ 3n ^{3} }\]

OpenStudy (erak):

yeah

OpenStudy (erak):

oh wait I see a way to simplify it more

jimthompson5910 (jim_thompson5910):

it's equivalent to the original expression but you can simplify that. So it's not the final answer

OpenStudy (erak):

yeah, those pesky n^2's. would that be the solution afterwards?

jimthompson5910 (jim_thompson5910):

\[\Large \Large \frac{ 2\sqrt{15n ^{3}} }{ 3n ^{3} } = \Large \frac{ 2\sqrt{n^2*15n} }{ 3n ^{3} }\] \[\Large \Large \frac{ 2\sqrt{15n ^{3}} }{ 3n ^{3} } = \Large \frac{ 2\sqrt{n^2}*\sqrt{15n} }{ 3n ^{3} }\] I'm sure you know what to do for the rest of the steps